OpenStudy (anonymous):

12% of what number is 66? please help! I just need the answer!

OpenStudy (king):

550

OpenStudy (saifoo.khan):

Five fifty?

OpenStudy (saifoo.khan):

wow.

OpenStudy (anonymous):

thx

OpenStudy (king):

ure welcome!!

OpenStudy (anonymous):

Let the number be x. 12% of x = 66 12/100 * x = 66 12x/100 = 66 12x = 6600 x = 6600/12 = 550